OpenStudy (anonymous):

4x(x-6)-5x(x-5)=-2 solve for x

OpenStudy (anonymous):

4x^2 - 24x - 5x^2 + 25x = -2 -x^2 +x = -2 -x^2 + x +2 = 0 x^2 -x -2 = 0 (x-2) (x+1) = 0 x = 2, -1

OpenStudy (anonymous):

is this right so far 4x^2-24-5x^2=25x=-2

OpenStudy (anonymous):

oops thats +25x

OpenStudy (anonymous):

wow i actually got one right lol